OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2.

OpenStudy (anonymous):

You can use the slope formula for "m":\[m = \frac{ y _{2} - y _{1} }{ x _{2} - x _{1} }\]where you are given points:\[(x _{1}, y _{1})\]and\[(x _{2}, y _{2})\]

OpenStudy (anonymous):

\[\frac{ -3 }{ 2 } = \frac{ 6 - (-3) }{ x - 10 }\]Can you solve it from here or do you need more help?

OpenStudy (anonymous):

I think I can handle it; thank you very much for the help! :)

OpenStudy (anonymous):

uw!